jhonyy9 (jhonyy9):

given p and q two primes greater than 2 P=2a+1 and q=2b+1 with a and b naturals prove that for n > 2 n=a+b+1 is always true - what proof method may be more usefully by contradiction or math.induction or ... ?

jhonyy9 (jhonyy9):

@ganeshie8 @mathmate @FaiqRaees @TheSmartOne @Directrix @Kainui

OpenStudy (drmoss):

@DrMoss :(

OpenStudy (loser66):

What is n?

jhonyy9 (jhonyy9):

n is natural greater than 2 wrote there

OpenStudy (loser66):

for all n?

jhonyy9 (jhonyy9):

yes for all n greater than 2

OpenStudy (loser66):

if p =3, q =5, then p =2*1 +1, q = 2*2 +1, hence a =1,b = 2 , then n = 1+2 +1 =4 That is the least of prime p, q >2, therefore with n=3, how can you have the proof valid?

OpenStudy (loser66):

n is always odd>3, that makes more sense. To me, p = 2a+1, then \(a=\dfrac{p-1}{2}\) q=2b+1, then \(b=\dfrac{q-1}{2}\) Let \(n = a+b+1=\dfrac{p-1}{2}+\dfrac{q-1}{2}+1=\dfrac{p-1+q-1+2}{2}=\dfrac{p+q}{2}\)

OpenStudy (mathmate):

If Goldbach's conjecture were accepted as true, it will take care of all the odd naturals >2. In other words, if you can prove the above, you have proved Goldbach's conjecture.

jhonyy9 (jhonyy9):

@Loser66 for n=3 this is valid really - look please there are p,q and n greater than 2 p=2a+1 and q=2b+1 so a = (p-1)/2 and b=(q-1)/2 - yes ? than n=a +b +1 so for p=3 =>a=3-1/2 =>a=1 for q=3 =>b=(3-1)/2 =>b=1 that we have n= 1+1+1 so n=3 do you believe it now that is true for n=3 ? hope helped

jhonyy9 (jhonyy9):

@ganeshie8 what is your opinion please about above wrote ? thank you

OpenStudy (zzr0ck3r):

I am confused about the question. Given primes p and q > 2 we know they are odd so p=2a+b, q=2b+1 Now we are saying that for all n>2 we have n=a+b+1? I am confused that it says "given p and q"

OpenStudy (mathmate):

Agree, the question by itself it's not clear. @jhonyy9 added the following comments. `n is natural greater than 2 wrote there ` `(for all n?)` `yes for all n greater than 2` It is not clear if it was part of the question, or if it was interpreted by OP. My previous response was based on his comments being valid.

jhonyy9 (jhonyy9):

@mathmate sorry what is not clearly there ? in the text of this question i wrote n>2 and n natural - so what is difficile in this - sorry but i dont understand it hope is clearly now for everybody - ty.

OpenStudy (mathmate):

What is not clear to me are the following: 1. As @loser66 says, `what is n?` If a and b are naturals, than a+b+1 is always natural. Is that what you needed to prove? or n>2 is what you need to prove? 2. You mentioned later that `yes for all n greater than 2`, is that your interpretation or is it part of the question? Why is it not mentioned in the question? 3. for the initial part of the question: `given p and q two primes greater than 2 ` is it for the two particular given p and q, or "given p and q are \(any\) primes greater than two" ? If the original question is in a different language, I suggest the original question be posted to avoid confusion. Translating mathematical text is very tricky.

OpenStudy (mathmate):

* `for **all** n?` `yes for all n greater than 2 `

jhonyy9 (jhonyy9):

@mathmate sorry but this may be more clearly than i ve wrote just for n > 2 and that always exist p and q ,p=2a+1 and q=2b+1 so p,q primes greater than 2 prove that this statement n= a +b +1 is true this may be more understandably text of above wrote exercise ? thank you